After bowling four games, Rita's average was 99. What must she score in the 5th game to average 100?
Answer by richwmiller(17219)

100-99=1
100-99=1
100-99=1
100-99=1
1+1+1+1=4
100+4=104 needed
(4*(99)+x)/5=100
(4*(99)+x)=500
396+x=500
x=104 needed